OpenStudy (anonymous):

Find the lowest common denominator that would eliminate the fractions in this equation. 3a/7+2/3=7/9

OpenStudy (anonymous):

So the answer is 36

OpenStudy (unklerhaukus):

seven times nine is sixty three try it and see what happens. \[\frac{3a}{7}+\frac{2}{3}=\frac{7}{9}\] multiply both sides by sixty three \[63\times\frac{3a}{7}+63\times\frac{2}{3}=63\times\frac{7}{9}\] \[9\times 3a+23\times 2=7 \times 7\] \[27a+46=49\]

OpenStudy (unklerhaukus):

\[27a=3\] \[a=\frac19\]